# Repetitive Contrastive Learning Enhances Mamba’s Selectivity in Time Series Prediction

### v1

@@ Abstract
journal: arxiv
@@
{highlights}
A token-level training paradigm named Repeating Sequence Augmentation is proposed. Through the repetition of timesteps and the introduction of noise, combined with intra-sequence and inter-sequence contrastive learning, the parameters of the Mamba block are enabled to identify key timesteps and ignore noise.
RCL is integrated into the training process of various Mamba-based models, with the parameters obtained from RCL being utilized as the initialization for the backbone model’s parameters, thereby further enhancing the temporal prediction capabilities of the backbone model.
It is demonstrated that RCL can significantly improve the performance of backbone models, and its broad effectiveness for Mamba-based models is proven without incurring additional memory overhead through experiments.
The impact of different parameter replacement methods and freezing techniques is experimentally analyzed.
Two metrics are proposed to measure the selective capabilities of Mamba. The effectiveness of RCL is demonstrated from theoretical, qualitative, and quantitative perspectives.
@@ 1 Introduction
Time series forecasting (TSF) has become indispensable across a range of critical domains, including financial markets (Li et al., 2023), traffic management (Cheng et al., 2023), electricity consumption prediction (Sun and Zhang, 2023), scientific computing (Cruz-Camacho et al., 2024), and weather forecasting (Zhang et al., 2022a). TSF leverages sequential data—often irregular, incomplete, or noisy—to predict future trends based on past observations. Yet, fully reliable forecasting remains elusive, largely due to the opaque generative processes behind time series data. These complexities are exacerbated by uneven sampling, missing or redundant entries, and unpredictable disturbances.(Zhu et al., 2023)(Ramponi et al., 2019) In such a setting, designing models that not only process temporal sequences but also learn to focus on the most informative and structurally meaningful parts of the data becomes vital—hinting at the importance of aligning training objectives with the intrinsic dynamics of time series itself.(Nam et al., 2024)
Deep learning has advanced significantly in the time series domain, with much of the attention placed on architectural innovations, particularly in transformer-based models (Wen et al., 2023). Despite their success in NLP, transformers often struggle with the sequential and noisy nature of time series, underperforming compared to traditional architectures like CNNs and MLPs (Zeng et al., 2022). These conventional models, while limited in modeling long-range dependencies, excel at handling noise and capturing local patterns, contributing to their relative success in this domain. The emergence of the Mamba model (Gu and Dao, 2024) has further shifted attention toward architectures tailored for temporal data. Leveraging a selective state-space mechanism (Huang et al., 2024; Li et al., 2024), Mamba addresses the quadratic complexity of transformers while preserving their long-range propagation ability. Its success in recent time series forecasting models, such as TimeMachine (Ahamed and Cheng, 2024) and Bi-Mamba (Liang et al., 2024), underscores the potential of this approach. However, these efforts often overlook a core challenge in time series: enabling models to identify and prioritize key timesteps amidst noise, a capability essential for robust and interpretable forecasting.
@@
Although Mamba introduces a degree of selectivity by generating its state transition and input matrices dynamically based on time steps, this very mechanism also makes it more sensitive to temporal fluctuations and noise. As illustrated in Figure 1, Mamba exhibits two key limitations when predicting time series. On one hand, it fails to effectively focus on salient time steps, resulting in an inability to fit extreme cases such as sharp declines or rapid rises. On the other hand, Mamba is highly sensitive to noise, where even minor noise disturbances can lead to significant deviations in prediction results. This is attributed to insufficient selective capabilities of Mamba, causing noise to accumulate progressively during sequence modeling, ultimately amplifying small noise into large deviations. These limitations arise from its origins in natural language processing (NLP), where each token typically carries rich semantic meaning. In contrast, time series data often exhibit irregular sampling, low signal-to-noise ratios, and a lack of contextual semantics, making them fundamentally different from NLP data. Prior studies have also shown that directly applying NLP-inspired architectures to time series tasks yields suboptimal results (Zhang et al., 2024).
@@ 2.2 Mamba Block
The Mamba block, (Gu and Dao, 2024), consists of two parts : selection and State Space Model (SSM), as shown in Fig. 2. Firstly, the input X undergoes a one-dimensional convolution (Conv1d) to extract local features, followed by Linear Projection that maps it to matrices B, C, and $\Delta$.
[…]
@@ B $=\text{fc}(\textbf{X}_{c}),\quad\textbf{C}=\text{fc}(\textbf{X}_{c})$
Finally, Mamba inputs $\overline{\textbf{A}}$, $\overline{\textbf{B}}$, C, $\Delta$, and X into the SSM, and uses residual connections.
$\textbf{H}=\text{SSM}(\overline{\textbf{A}},\overline{\textbf{B}},\textbf{X})\cdot\sigma(\text{fc}(\textbf{X}))$ (3)
where fc is fully connected layers, and $\sigma$ is SiLU activation function. The computational process of the State Space Model (SSM) can be succinctly represented as follows:
@@
Figure 2: The structure of the mamba block.
@@ 2.3 Definition of Mamba’s Selectivity
The selectivity of Mamba primarily stems from its unique Selective State Space Model (SSM). The computational process of the SSM is shown in Eq.4. Unlike traditional SSMs (Gu et al., 2022), the state transition matrix $\overline{\textbf{A}}$ and the input matrix $\overline{\textbf{B}}$ are derived from the current timestep. As a result, the state transition matrix and input matrix generated based on the timestep can selectively decide whether to retain more historical state information or incorporate more information from the current timestep (Gu and Dao, 2024).
@@ 2.4 Definition of a Selectivity Measurement
Since Mamba, unlike Transformer, does not provide explicit attention scores to intuitively measure selectivity, understanding how it retains or discards information requires alternative strategies. While Mamba is based on a Selective State Space Model (SSM) rather than a gated recurrent mechanism, it shares with RNNs and LSTMs the key characteristic of processing sequences in a token-by-token manner, rather than consuming entire sequences simultaneously as Transformers do. This temporal nature of computation motivates us to draw inspiration from prior works analyzing memory and information retention in RNNs, where token-level dynamics — such as the evolution of hidden states across time — have been used to study memory behaviors (Zhang et al., 2020; Haviv et al., 2019).
Building on this perspective, we propose two quantitative metrics to assess the selectivity of Mamba. First, based on Eq. 4, we compute the correlation between the current hidden state $h_{t}$, the previous hidden state $h_{t-1}$, and the current input $x_{t}$, normalizing the results to sum to 1. We define the correlation with $x_{t}$ as the memory score $s_{t}$ at the current time step. Following the definitions of memory and ignoring from Section 2.3, we categorize time steps with $s_{t}>0.7$ as Significant Memory (SM), those with $s_{t}<0.3$ as Significant Ignoring (SI), and the remainder as Normal (NR). This approach is in line with previous efforts to quantify memory at each time step in sequential models by examining how hidden representations evolve over time (Ming et al., 2017). Based on these memory scores, we define two Selectivity Measurements: Focus Ratio and Memory Entropy.
@@ 3.1 Repeating Sequence Augmentation
[…]
$,\displaystyle\textbf{X}_{\text{rep}}=\{\textbf{X}_{1,1},...,\textbf{X}_{1,n_{t}},$$...,\textbf{X}_{i,1},...,$
$\textbf{X}_{i,n_{t}},...,\textbf{X}_{s,1},...,\textbf{X}_{s,n_{t}}\}$
@@
$\textbf{X}_{\text{aug},i}$ =$=\{\textbf{X}_{i,1},\hat{\textbf{X}}_{i,2},\hat{\textbf{X}}_{i,3}\}$
@@
where $\text{Noise}_{\alpha}$ and $\text{Noise}_{\beta}$ represent weak and strong Gaussian noise, controlled by the variances $\sigma_{\alpha}$ and $\sigma_{\beta}$, ||||| | denotes the sequential concatenation of sequences. Since the impact of noise accumulates progressively during sequence modeling, gradually increasing the noise effectively amplifies the distance between time steps. As a result, the repeated time steps form a sequence of denoising targets with progressively increasing difficulty.
@@ 3.2 Repetitive Contrastive Learning
Intra-sequence contrast We hypothesize that if the Mamba Block possesses strong sequence selection capabilities, then the outputs $\{\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{i\cdot n_{t}+1,\text{aug}\}},\ldots,\textbf{H}_{\{i\cdot n_{t}+n_{t}-1,\text{aug}\}}\}$ of the augmented sequence at the same time step should exhibit high similarity, while ignoring progressively increasing noise. Conversely, the outputs $\textbf{H}_{\{i\cdot n_{t},\text{aug}\}}$ at the current time step and $\textbf{H}_{\{(i+1)\cdot n_{t},\text{aug}\}}$ at the subsequent time step should have low similarity. Therefore, we define outputs at the same time step as positive examples, while outputs at the current and subsequent time steps serve as negative examples. The objective is to minimize the distance between positive examples and maximize the distance between negative examples within the sequence, thereby enhancing the Mamba Block’s sequence selection capabilities. Specifically, we use $\textbf{H}_{\{i\cdot n_{t},\text{aug}\}}$ as an anchor to form $n_{t}-1$ positive samples and one negative sample, measuring similarity between samples using cosine similarity and employing the InfoNCE loss function Oord et al. (2018).
$\mathcal{L}_{\text{Intra}}=$ $-\frac{1}{s-1}\sum_{i=0}^{s-2}\frac{1}{n_{t}-1}\sum_{z=1}^{n_{t}-1}\log\frac{\exp(\text{sim}(\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)}{\exp(\text{sim}(\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{(i+1)\cdot n_{t},\text{aug}\}})/\tau)}$ (8)
@@
$\mathcal{L}_{\text{Inter}}=-\frac{1}{s-1}$ $\sum_{i=0}^{s-2}\frac{1}{n_{t}}\sum_{z=0}^{n_{t}-1}\log\frac{\exp(\text{sim}(\textbf{H}_{i},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)}{\exp(\text{sim}(\textbf{H}_{i},\textbf{H}_{i+1})/\tau)}$ (10)
@@ 4 Experiment
We conducted extensive experiments to validate the effectiveness of our method. In Section 4.2, we compare the prediction performance of various Mamba-based models—Mamba (Gu and Dao, 2024), iMamba, TimeMachine (Ahamed and Cheng, 2024), Bi-Mamba (Liang et al., 2024) and SiMBA(Patro and Agneeswaran, 2024)—both with and without pre-trained parameters across multiple datasets: ETTh1, ETTh2, ETTm1, ETTm2, Traffic, and Electricity.
In Section 4.1, we provide an overview of the experimental setup and basic information. In Section 4.2, we demonstrate the performance improvement of the backbone model achieved by RCL. In Section 4.3, we compare RCL with other pretraining methods. In Section 4.4, we present the ablation study results of RCL. In Section 4.5, we analyze the impact of RCL parameters. In Section 4.6, we compare RCL with other temporal prediction models and achieve state-of-the-art results. In Section 4.7, we examine the effects of different parameter substitution ratios and freezing methods. Additionally, in B, we list additional experimental results.
@@ 4.1.1 Mamba-based Baseline
    Mamba (Gu and Dao, 2024): Mamba is a new Selective State Spaces model proposed by Albert Gu and Tri Dao in 2024.(Li et al., 2024) It demonstrates outstanding performance in sequence modeling through its selective state space formulation, effectively capturing long-range dependencies while maintaining computational efficiency.
@@
    TimeMachine (Ahamed and Cheng, 2024): TimeMachine, introduced by Md Atik Ahamed and Qiang Cheng in 2024, is designed for long-term sequence forecasting. By integrating channel-independent and channel-mixed modeling approaches, it achieves state-of-the-art performance. The architecture incorporates four Mamba blocks, optimizing predictive capability over extended sequences.
@@
    Bi-Mamba (Liang et al., 2024): Bi-Mamba was proposed in 2024, Bi-Mamba extends the Mamba framework by adaptively capturing both internal and inter-series dependencies in multivariate time series data. The model introduces forget gates, enabling it to retain relevant historical information over extended time periods, thereby enhancing its forecasting accuracy.
@@
    SiMBA (Patro and Agneeswaran, 2024): SiMBA is a hybrid architecture that combines Mamba-based sequence modeling with EinFFT, a novel FFT-based channel mixer. It is designed to overcome Mamba’s instability when scaling, offering a stable and efficient solution for large-scale sequence tasks. SiMBA refines selective state space models for improved scalability and performance in visual recognition. Due to space limitations, we omitted experiments whose output length longer than 96 for the SiMBA model.

### v2

@@ Abstract
journal: Neural Networksequalequalfootnotetext: Same Contribution
@@ 1 Introduction
Time series forecasting (TSF) has become indispensable across critical domains such as financial markets (li2023mastermarketguidedstocktransformer), traffic management (cheng2023deeptransportlearningspatialtemporaldependency), electricity consumption (Sun2023), scientific computing (10.1145/3682060), and weather prediction (ZHANG2022111768). TSF leverages sequential data—often irregular, incomplete, or noisy—to predict future trends based on past observations. Deep learning has advanced the field substantially, with increasing attention on architectural innovations, particularly transformer-based models (wen2023transformerstimeseriessurvey). Yet, conventional models like CNNs and MLPs remain competitive due to their robustness against noise and ability to capture local patterns (zeng2022transformerseffectivetimeseries). More recently, the Mamba model (gu2024mambalineartimesequencemodeling), powered by selective state-space mechanisms (huang2024mlmambaefficientmultimodallarge; li2024mambandselectivestatespace), has gained traction, inspiring variants such as TimeMachine (ahamed2024timemachinetimeseriesworth) and Bi-Mamba (liang2024bimambabidirectionalmambatime).
Despite these advances, fully reliable forecasting remains elusive. The challenges stem from the opaque generative processes of time series data, exacerbated by uneven sampling, missing or redundant entries, and unpredictable disturbances (zhu2023networkedtimeseriesprediction; ramponi2019tcganconditionalgenerativeadversarial). Transformer architectures, though powerful in NLP, often underperform in these settings, struggling with sequential irregularity and noise. While state-space approaches like Mamba alleviate complexity and preserve long-range dependencies, they often overlook a core challenge: enabling models to identify and prioritize the most informative timesteps. Addressing this issue is vital for robust and interpretable forecasting (nam2024adversariallearningapproachirregular). In particular, as we will discuss next, Mamba’s selective mechanism—though promising—introduces its own limitations when applied to time series forecasting.
@@
Although Mamba introduces a degree of selectivity by generating its state transition and input matrices dynamically based on time steps, this very mechanism also makes it more sensitive to temporal fluctuations and noise. As illustrated in Figure 1, Mamba exhibits two key limitations when predicting time series. On one hand, it fails to effectively focus on salient time steps, resulting in an inability to fit extreme cases such as sharp declines or rapid rises. On the other hand, Mamba is highly sensitive to noise, where even minor noise disturbances can lead to significant deviations in prediction results. This is attributed to insufficient selective capabilities of Mamba, causing noise to accumulate progressively during sequence modeling, ultimately amplifying small noise into large deviations. These limitations arise from its origins in natural language processing (NLP), where each token typically carries rich semantic meaning. In contrast, time series data often exhibit irregular sampling, low signal-to-noise ratios, and a lack of contextual semantics, making them fundamentally different from NLP data. Prior studies have also shown that directly applying NLP-inspired architectures to time series tasks yields suboptimal results (zhang2024integrationmambatransformer).
@@ 2.2 Mamba Block
The Mamba block, (gu2024mambalineartimesequencemodeling), consists of two parts : selection and State Space Model (SSM), as shown in Fig. 2(a). Firstly, the input X undergoes a one-dimensional convolution (Conv1d) to extract local features, followed by Linear Projection that maps it to matrices B, C, and $\Delta$.
$\textbf{X}_{c}$ $=\text{fc}(\sigma(\text{Conv1d}(\textbf{X})))$ (1)
@@ B $=\text{fc}(\textbf{X}_{c}),\quad\textbf{C}=\text{fc}(\textbf{X}_{c})$
Finally, Mamba inputs $\overline{\textbf{A}}$, $\overline{\textbf{B}}$, C, $\Delta$, and X into the SSM with residual connections and passes the output through a Linear Projection.
$\textbf{H}=\text{fc}(\text{SSM}(\overline{\textbf{A}},\overline{\textbf{B}},\textbf{X})\cdot\sigma(\text{fc}(\textbf{X})))$ (3)
where fc is fully connected layers, and $\sigma$ is SiLU activation function. The process of sequence modeling with Mamba is illustrated in Fig. 2(b). The computational process of the State Space Model (SSM) can be succinctly represented as follows:
@@
Figure 2: (a)The structure of the mamba block. (b)The Process of Sequence Modeling with Mamba
@@ 2.3 Definition of Mamba’s Selectivity
The selectivity of Mamba primarily stems from its unique Selective State Space Model (SSM). The computational process of the SSM is shown in Eq.4. Unlike traditional SSMs (gu2022efficientlymodelinglongsequences), the state transition matrix $\overline{\textbf{A}}$ and the input matrix $\overline{\textbf{B}}$ are derived from the current timestep. As a result, the state transition matrix and input matrix generated based on the timestep can selectively decide whether to retain more historical state information or incorporate more information from the current timestep (gu2024mambalineartimesequencemodeling).
@@ 2.4 Definition of a Selectivity Measurement
Since Mamba, unlike Transformer, does not provide explicit attention scores to intuitively measure selectivity, understanding how it retains or discards information requires alternative strategies. While Mamba is based on a Selective State Space Model (SSM) rather than a gated recurrent mechanism, it shares with RNNs and LSTMs the key characteristic of processing sequences in a token-by-token manner, rather than consuming entire sequences simultaneously as Transformers do. This temporal nature of computation motivates us to draw inspiration from prior works analyzing memory and information retention in RNNs, where token-level dynamics — such as the evolution of hidden states across time — have been used to study memory behaviors (zhang2020assessingmemoryabilityrecurrent; haviv2019understandingcontrollingmemoryrecurrent).
Building on this perspective, we propose two quantitative metrics to assess the selectivity of Mamba. First, based on Eq. 4, we compute the correlation between the current hidden state $h_{t}$, the previous hidden state $h_{t-1}$, and the current input $x_{t}$, normalizing the results to sum to 1. We define the correlation with $x_{t}$ as the memory score $s_{t}$ at the current time step. Following the definitions of memory and ignoring from Section 2.3, we categorize time steps with $s_{t}>0.7$ as Significant Memory (SM), those with $s_{t}<0.3$ as Significant Ignoring (SI), and the remainder as Normal (NR). This approach is in line with previous efforts to quantify memory at each time step in sequential models by examining how hidden representations evolve over time (8585721). Based on these memory scores, we define two Selectivity Measurements: Focus Ratio and Memory Entropy.
@@ 3.1 Repeating Sequence Augmentation
$\textbf{X}_{i}\xrightarrow{\text{repeat}}$ $\ \textbf{X}_{i,1},.,\textbf{X}_{i,n_{t}}$ (6)
$\textbf{X}_{\text{rep}}=\{\textbf{X}_{1,1},.,\textbf{X}_{1,n_{t}},$ $.,\textbf{X}_{i,1},$.,
$\textbf{X}_{i,n_{t}},.,\textbf{X}_{s,1},.,\textbf{X}_{s,n_{t}}\}$
@@
$\textbf{X}_{\text{aug},i}$ $=\{\textbf{X}_{i,1},\hat{\textbf{X}}_{i,2},\hat{\textbf{X}}_{i,3}\}$
@@
where $\text{Noise}_{\alpha}$ and $\text{Noise}_{\beta}$ represent weak and strong Gaussian noise, controlled by the variances $\sigma_{\alpha}$ and $\sigma_{\beta}$, $||$ denotes the sequential concatenation of sequences. Since the impact of noise accumulates progressively during sequence modeling, gradually increasing the noise effectively amplifies the distance between time steps. As a result, the repeated time steps form a sequence of denoising targets with progressively increasing difficulty.
@@ 3.2 Repetitive Contrastive Learning
Intra-sequence contrast We hypothesize that if the Mamba Block possesses strong sequence selection capabilities, then the outputs $\{\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{i\cdot n_{t}+1,\text{aug}\}},\ldots,\textbf{H}_{\{i\cdot n_{t}+n_{t}-1,\text{aug}\}}\}$ of the augmented sequence at the same time step should exhibit high similarity, while ignoring progressively increasing noise. Conversely, the outputs $\textbf{H}_{\{i\cdot n_{t},\text{aug}\}}$ at the current time step and $\textbf{H}_{\{(i+1)\cdot n_{t},\text{aug}\}}$ at the subsequent time step should have low similarity. Therefore, we define outputs at the same time step as positive examples, while outputs at the current and subsequent time steps serve as negative examples. The objective is to minimize the distance between positive examples and maximize the distance between negative examples within the sequence, thereby enhancing the Mamba Block’s sequence selection capabilities. Specifically, we use $\textbf{H}_{\{i\cdot n_{t},\text{aug}\}}$ as an anchor to form $n_{t}-1$ positive samples and one negative sample, measuring similarity between samples using cosine similarity and employing the InfoNCE loss function oord2018representation.
$\mathcal{L}_{\text{Intra}}=$ $-\frac{1}{s-1}\sum_{i=0}^{s-2}\log\frac{\sum_{z=1}^{n_{t}-1}\exp(\text{sim}(\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)}{\sum_{z=1}^{n_{t}-1}\exp(\text{sim}(\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)+\exp(\text{sim}(\textbf{H}_{\{i\cdot n_{t},\text{aug}\}},\textbf{H}_{\{(i+1)\cdot n_{t},\text{aug}\}})/\tau)}$ (8)
@@
$\mathcal{L}_{\text{Inter}}=-\frac{1}{s-1}$ $\sum_{i=0}^{s-2}\log\frac{\sum_{z=0}^{n_{t}-1}\exp(\text{sim}(\textbf{H}_{i},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)}{\sum_{z=0}^{n_{t}-1}\exp(\text{sim}(\textbf{H}_{i},\textbf{H}_{\{i\cdot n_{t}+z,\text{aug}\}})/\tau)+\exp(\text{sim}(\textbf{H}_{i},\textbf{H}_{i+1})/\tau)}$ (10)
@@ 4 Experiment
We conducted extensive experiments to validate the effectiveness of our method. In Section 4.2, we compare the prediction performance of various Mamba-based models—Mamba (gu2024mambalineartimesequencemodeling), iMamba, TimeMachine (ahamed2024timemachinetimeseriesworth), Bi-Mamba (liang2024bimambabidirectionalmambatime) and SiMBA(patro2024simbasimplifiedmambabasedarchitecture)—both with and without pre-trained parameters across multiple datasets: ETTh1, ETTh2, ETTm1, ETTm2, Traffic, and Electricity.
In Section 4.1, we provide an overview of the experimental setup and basic information. In Section 4.2, we demonstrate the performance improvement of the backbone model achieved by RCL. In Section 4.3, we compare RCL with other pretraining methods. In Section 4.4, we present the ablation study results of RCL. In Section 4.5, we analyze the impact of RCL parameters. In Section 4.6, we compare RCL with other temporal prediction models and achieve state-of-the-art results. In Section 4.7, we examine the effects of different parameter substitution ratios and freezing methods. Additionally, in D, we list additional experimental results.
@@ 4.1.1 Mamba-based Baseline
    Mamba (gu2024mambalineartimesequencemodeling): Mamba is a new Selective State Spaces model proposed by Albert Gu and Tri Dao in 2024.(li2024mambandselectivestatespace) It demonstrates outstanding performance in sequence modeling through its selective state space formulation, effectively capturing long-range dependencies while maintaining computational efficiency.
@@
    TimeMachine (ahamed2024timemachinetimeseriesworth): TimeMachine, introduced by Md Atik Ahamed and Qiang Cheng in 2024, is designed for long-term sequence forecasting. By integrating channel-independent and channel-mixed modeling approaches, it achieves state-of-the-art performance. The architecture incorporates four Mamba blocks, optimizing predictive capability over extended sequences.
@@
    Bi-Mamba (liang2024bimambabidirectionalmambatime): Bi-Mamba was proposed in 2024, Bi-Mamba extends the Mamba framework by adaptively capturing both internal and inter-series dependencies in multivariate time series data. The model introduces forget gates, enabling it to retain relevant historical information over extended time periods, thereby enhancing its forecasting accuracy.
@@
    SiMBA (patro2024simbasimplifiedmambabasedarchitecture): SiMBA is a hybrid architecture that combines Mamba-based sequence modeling with EinFFT, a novel FFT-based channel mixer. It is designed to overcome Mamba’s instability when scaling, offering a stable and efficient solution for large-scale sequence tasks. SiMBA refines selective state space models for improved scalability and performance in visual recognition. Due to space limitations, we omitted experiments whose output length longer than 96 for the SiMBA model.